OpenStudy (anonymous):

Write the slope-intercept form of the equation of each line through the give point with the given slope through: (5,4) slope = 6/5

OpenStudy (anonymous):

use formula y-y1=m(x-x1)

OpenStudy (anonymous):

y-4=6/5 (x-5) 5y-4=6x-30 5y=6x-26

OpenStudy (anonymous):

thank same one (-2,-4) slope 1/2

OpenStudy (anonymous):

i gave you the formula try it yourself!!

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-4=\frac{6}{5}(x-5)\] \[y-4=\frac{6}{5}x-6\] \[y=\frac{6}{5}x-2\]

OpenStudy (anonymous):

there is a slight mistake in Sarkar's answer

OpenStudy (anonymous):

if you want to multiply both sides by 5 you have to multiply everything by 5

OpenStudy (anonymous):

i have erred 5y-20=6x-30 and not 4 thanks @satellite

OpenStudy (anonymous):

i would leave it at \[y=mx+b\] don't be scared of the fraction

OpenStudy (anonymous):

i get told off on school for leaving answers as fraction so i have a tendency to evade it

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[m=\frac{1}{2},x_1=-2,y_1=-4\] \[y-(-4)=\frac{1}{2}(x-(-2))\] \[y+4=\frac{1}{2}(x+2)\] \[y+4=\frac{1}{2}x+1\] \[y=\frac{1}{2}x-3\] steps always the same, a) put in point - slope formula b) distribute on the right c) solve for y

OpenStudy (anonymous):

c solve for y

OpenStudy (aravindg):

hmm ...i am late

OpenStudy (anonymous):

okay